# LLMs, You Can Evaluate It! Design of Multi-perspective Evaluation for Reports in Security Operation Centers

### v2

@@ 15 Low / 2 SOC Analyst, Network Engineer IT, Factory
We conducted semi-structured interviews with SOC personnel responsible for drafting and reviewing analysis reports in order to develop a checklist [rose2022co] that reflects SOC practitioners’ concerns about analysis reports and the operational context in which they are used. The purpose of these interviews was to elicit the implicit evaluation criteria that analysts rely on when judging the “quality” of a report and to make those criteria explicit as checklist items, focusing on the perspectives deemed particularly important. For the “report elements that should be documented” identified earlier, generic criteria and metrics for text quality can be applied to some extent. In this study, however, we deliberately focused on the perspectives that SOC practitioners themselves strongly recognize as problematic in their daily work, and prioritized clarifying which aspects should be checked most carefully. For example, while purely formal aspects such as spelling and typographical errors cannot be entirely ignored, it is more likely that, in analysis reports, practical “quality” is determined by content-related factors such as the technical accuracy of the detected event, the adequacy of the description of the operational context, and the concreteness and feasibility of the recommended response actions. We therefore considered it essential to capture these viewpoints appropriately and to ensure that they are explicitly and transparently represented in the checklist.
@@ 3.3.1 Study Design
We conducted interviews with 15 participants from 8 different SOC-related organizations. The participants included senior analysts responsible for alert analysis and customer-facing reports, team leaders, and managers, all of whom are routinely involved in at least one part of the report lifecycle—authoring, reviewing, or approving reports. Participants were recruited via a B2B interview platform, from within the authors’ own organization, and from partner companies. The participants included members of SOCs responsible not only for IT but also for OT/IoT environments, thereby ensuring diversity from multiple perspectives. Detailed demographic and organizational attributes of the participants are provided in the Table 1. All interviews were conducted online using conferencing tools such as Microsoft Teams. Wherever feasible, interviews were conducted by multiple authors to mitigate potential biases in questioning and interpretation. With participants’ consent, all interviews were audio-recorded and subsequently transcribed, yielding a total of 337 coded comments. Prior to analysis, the transcripts were anonymized and stored with appropriate security safeguards to ensure the ethical handling of the data. For participants from organizations external to the authors’ affiliation, we provided an honorarium set at a uniform amount across all participants. To allow the conversation to develop naturally while maintaining coverage of key topics, we adopted a semi-structured interview format, with each session lasting approximately 30 to 60 minutes. Example interview questions are listed in the Table 2.
@@ 1-10. How long does it take to create a single report?
Preliminary User Study 2-1. Are there any tools or automation systems you use to make report creation more efficient? If so, what are they?
@@ 2-16. What aspects of reporting do you find particularly challenging?
As our analysis method, we employed template analysis (TA), which begins with a set of a priori themes of interest and iteratively refines and extends the coding template by incorporating newly emerging themes. TA is particularly useful in domains such as SOC research, where prior work is still limited and the researcher has only a partial understanding of the concepts that need to be identified in the data. [bushra202299False] First, we conducted an initial round of coding based on the descriptions obtained from the literature review and developed a preliminary set of thematic codes. We then analyzed the interview data, and whenever we encountered noteworthy segments that did not match any existing thematic code, we either introduced new codes or refined and extended the existing themes. Coding was performed iteratively three times, with intermediate code sets reviewed by SOC domain experts, and the final coding scheme was established based on the outcomes of these iterations. Moreover, as no new codes emerged after the 14th interview, we judged that thematic saturation had been reached and therefore did not conduct additional interviews beyond the 15 participants. In the following, we describe the evaluation criteria for assessing report quality that were derived from the interview results.
@@ 3.3.2 Implicit Evaluation Criteria for Security Report Quality
This section organizes the implicit criteria that SOC practitioners use when assessing the quality of security reports and refines them into explicit dimensions for evaluating report quality. Although most of the organizations we interviewed did not maintain formal documentation explicitly defining “report evaluation criteria,” we found that experienced Tier-2–level analysts, recipient-side managers, and SIRT members effectively applied similar criteria when reviewing reports. For an overview of interview findings other than the report evaluation criteria, refer to Table 3. Moreover, most comments pointed to shortcomings in the substantive content of the reports rather than in their formal qualities. In the following, we present the resulting evaluation criteria and related comments, organized according to the three categories of items that should be included in the report.
@@
Criterion (1): Clarity and persuasiveness of the bottom-line conclusion for rapid situation understanding. Respondents emphasized that analysis reports should present a clear and persuasive bottom-line conclusion that enables recipients to rapidly grasp the situation and make informed decisions. In particular, reports are expected to explicitly state whether the alert is valid, its urgency, and the recommended stance toward the situation, together with the reasoning that supports these judgments. However, participants noted that many reports merely enumerate factual observations without articulating a decisive conclusion. “High-quality reports clearly state whether an alert should be treated as a true positive or not, how urgent it is, and why that judgment was made. Without that bottom line, it is hard to decide the next step quickly.” –- UserID 1. “For decision-making, the key is whether the report clearly explains what the problem is, why it matters, and what conclusion we should draw from the analysis.” –- UserID 3.
@@
Criterion (2): Quality of concreteness and prioritization in the proposed next actions. Participants stressed that reports should not stop at analysis results but should concretely describe the next actions to be taken, along with their relative priority. Effective reports distinguish between actions that must be taken immediately and those that are optional or conditional, enabling recipients to allocate resources appropriately. In contrast, reports that simply list findings without actionable guidance were considered insufficient for operational use. “It is important to clearly state what should be done next and to indicate priorities, for example, whether an action is mandatory or should be handled if resources allow.” –- UserID 9. “Readers need to understand not only what issues exist, but also what actions they themselves are expected to take in response to those issues.” –- UserID 6.
@@
Criterion (3): Adequacy and clarity of the described customer or business impact. Interviewees regarded it as essential that reports clearly describe how the alert affects the customer’s business, operations, or productivity, from the customer’s perspective. Reports that focus solely on technical details while leaving business impact ambiguous make it difficult for recipients to judge acceptability or risk. When possible, quantifying the impact was seen as particularly valuable for facilitating coordination and decision-making. “If the impact on business or operations is clearly described or even roughly quantified, it becomes much easier to explain the situation and move forward.” –- UserID 11. “Ultimately, customers want to know what caused the issue and whether the situation is truly acceptable. If that is unclear, the report leaves them uneasy.” –- UserID 14.
@@
Criterion (4): Appropriateness of tailoring the content, tone, and level of detail to the recipient’s role and level of understanding. Respondents highlighted the importance of adjusting the content, tone, and level of technical detail according to the recipient’s role and expertise, in order to reduce cognitive burden and align perceptions of urgency. Participants pointed out that reports written in a uniform style often fail to bridge the gap between frontline analysts and decision-makers, resulting in misunderstandings or delayed responses. “Readability is critical. Each recipient has a different level of understanding, so the explanation and terminology need to be adjusted accordingly.” –- UserID 2. “What matters most is narrowing the gap in sense of urgency between the frontline and decision-makers. Reports should reflect the context and discussions that led up to the analysis.” –- UserID 10.
@@
Evaluation criterion (5): Accuracy and explanatory strength of factual descriptions, including explicit assumptions and constraints Analysts emphasized that reports must accurately describe what occurred during the event and provide explanations that convincingly justify the conclusions drawn. Beyond factual correctness, respondents highlighted the importance of explicitly stating the assumptions, analytical grounds, and constraints underlying the analysis. Reports that omit these elements were considered difficult to trust, as readers cannot assess whether the conclusions are well-founded. “The most important thing is to accurately describe what actually happened and to explain it in a way that the reader can reasonably accept.” –- UserID 7. “If the assumptions, grounds, or constraints behind the analysis are unclear, it becomes hard to judge whether the situation has really been understood correctly.” –- UserID 4.
@@
Evaluation criterion (6): Appropriateness and depth of on-site contextual information usage, such as asset roles and operational background. Participants reported that proper interpretation of alerts requires the use of on-site contextual information, including asset roles, system usage, and their position within operational or business processes. Without such context, analysts often cannot determine whether an alert represents a genuine issue or a false positive. Respondents noted that insufficient use of operational context frequently leads to overly conservative or ambiguous conclusions. “Operational context is indispensable. Depending on the role of the asset and how it is used, the priority and meaning of the alert can change completely.” –- UserID 13. “If information such as asset roles, whether the traffic is business-related, or results of on-site interviews is missing, we often cannot clearly determine whether an alert is a false positive.” –- UserID 5.
@@
Evaluation criterion (7): Adequacy of technical depth in describing the event beyond simple enumeration of logs. Rather than merely listing communication records or log entries, reports are expected to provide technically meaningful explanations of the event, such as which systems communicated, for what purpose, and how that behavior should be interpreted. Several participants expressed frustration that the burden of interpreting raw technical data is often left entirely to the recipient. “SOC reports often just list IP addresses that communicated, and we have to re-investigate where those systems are and what they are used for.” –- UserID 6. “If the report also explained the meaning of the communication, it would significantly reduce the investigation workload on our side.” –- UserID 11.
@@
Evaluation criterion (8): Robustness of anomaly identification through multiple analytical approaches, such as trends and comparisons. Finally, interviewees underscored the importance of identifying anomalies using multiple analytical perspectives, including temporal trends, comparisons with historical data, and correlations across different logs. Reports that rely solely on isolated log entries were considered insufficient for understanding whether observed behavior is truly abnormal. Participants noted that temporal and comparative analyses are still underutilized in many current reports. “What we really want to know is what has changed compared to the past, but that perspective is often missing in reports.” –- UserID 9. “Showing insights derived from differences in trends or combinations of logs would make anomalies much clearer.” –- UserID 3.
@@
Evaluation criterion (10): Clarity and verifiability of the analysis process through explicit and coherent links between evidence and conclusions. It was considered critical that reports clearly document the reasoning process by explicitly linking evidence to conclusions, so that a third party could trace and verify how judgments were reached. Participants pointed out that reports often fail to sufficiently document how alternative explanations, such as false positives, were examined, which undermines confidence in the conclusions and the overall quality of the analysis. “It is important that there are no logical gaps or contradictions in how the analysis results lead to the conclusions, and that a third party could reproduce the same reasoning.” –- UserID 7. “Especially for determining whether an alert is a false positive, the evidence and reasoning must be clearly connected. If that part is weak, the quality of the entire report appears low.” –- UserID 11.
@@ 3.4 Development of the Analyst-wise Checklist Draft
This checklist was developed by systematically integrating 1) the structural elements that should be included in analysis reports, extracted through a literature review, with 2) the implicit evaluation criteria used in practice by SOC analysts, elicited through interviews. The integrated item set was then finalized through expert review and application tests. This approach is also widely known as a general method for constructing evaluation checklists [hales2008development].
@@
When designing checklist items, our primary requirement is that they remain actionable in day-to-day SOC operations. Simply turning interviewees’ concerns into items risks creating demands that exceed the authority or resources of a typical SOC organization[rose2022co]. We therefore introduced a step that distills practically feasible elements from the interview findings and refines them into checklist items. For example, for evaluation criterion (6): use of on-site contextual information, interviewees repeatedly reported that the lack of local context often prevents them from deciding whether an alert is a false positive. However, it is unrealistic for a standard SOC to maintain detailed, up-to-date business context for every customer environment. Accordingly, our checklist does not merely ask whether incidents are analyzed with customer context in mind, but also whether analysts explicitly state which contextual information was available and organize missing yet decision-critical information as concrete confirmation requests. This yields a realistic checklist that assumes a division of roles with on-site staff while still reducing ambiguity caused by insufficient context.
The Analyst-wise Checklist is also applicable to general tasks by human in SOCs because of the following concept: 1) the granularity of checklist items are generalized for the use in multiple domains, 2) for operational difficulty and flexibility in real-world systems, specific evaluation criteria are removed, 3) instead of applying the entire checklist, only relevant items for each analysis report are chosen.
@@ 3.5 Refinement through Expert Review and Evaluation Tests
As a final validation step, one of the three experts and one of the authors with more than seven years of SOC-related experience independently applied the checklist to ten real-world security alert analysis reports. For each checklist item, they rated the extent to which it was satisfied on a five-point scale and recorded the results. For items where their ratings diverged substantially in direction, e.g., one judged the item to be largely satisfied while the other judged it to be hardly satisfied, they reconciled their interpretations and refined the item descriptions and wording accordingly.
We confirmed that for more than 60% of the checklist items, the two raters agreed on the direction of judgment, i.e., whether the item was satisfied or not, and we then finalized the checklist. Table 4 shows a subset of the checklist with mappings to the evaluation criteria; The complete checklist and checklist item types are provided in Table 5,  6.

### v3

@@ 15 Low / 2 SOC Analyst, Network Engineer IT, Factory
We conducted semi-structured interviews with SOC practitioners responsible for drafting and reviewing analysis reports in order to develop the Analyst-wise Checklist. It can reflect SOC practitioners’ concerns about analysis reports and the operational context into the checklist [rose2022co]. More specifically, the purpose of these interviews is to elicit the implicit evaluation criteria that SOC practitioners use when judging the quality of analysis reports. One might think that generic criteria and metrics for text quality can be applied to the three categories and eleven items obtained in the previous process. However, we focus on the perspectives that SOC practitioners strongly recognize as problematic in their work, and prioritize clarifying which aspects should be checked most carefully. For example, while purely formal aspects such as spelling and typographical errors are reasonably important, it is more likely that the quality of analysis reports is determined by content-related factors, such as adequate descriptions of the operational context and feasible actions. It is therefore essential to capture these evaluation criteria appropriately in the checklist items.
@@ 3.3.1 Study Design
We conducted interviews with fifteen participants from eight different SOC-related organizations. Participants included veteran analysts, managers, and SIRT members, all of whom were involved in at least one stage of analysis reports, such as authoring, reviewing, and approving reports. The participants were recruited via an online interview platform from IT/OT/IoT environments, including the authors’ affiliations and partner companies. Detailed organizational attributes of the participants are shown in Table 1. All the interviews were performed on Microsoft Teams with multiple authors in order to mitigate potential biases in questions and their interpretation. Upon the participants’ consent, all the interviews were audio-recorded and subsequently transcribed, yielding a total of 337 comments. Prior to analysis, the transcripts were anonymized and stored with appropriate security safeguards for research ethics. For the participants from organizations outside the authors’ affiliations, we provided a reward of a common amount for all the participants. To encourage discussion based on the conversation of interviews, we adopted a semi-structured interview format with approximately thirty to sixty minutes. An actual template of interview questions are shown in Table 2.
@@ 1-10. How long does it take to create a single report?
User Study 2-1. Are there any tools or automation systems you use to make report creation more efficient? If so, what are they?
@@ 2-16. What aspects of reporting do you find particularly challenging?
We employ template analysis, which begins with a prior set of themes with interests and iteratively refines them by extending templates by incorporating new themes. It is considered that template analysis is useful in domains such as SOC research when prior work has a partial understanding of the concepts that need to be identified in advance [bushra202299False]. First, we conducted an initial round of the coding process based on the descriptions obtained from the literature review and developed an initial set of thematic codes. We then analyzed the interview data and, whenever we found new codes that did not match any existing codes, we refined the existing themes by incorporating the new codes. The above coding process was performed iteratively three times, with code reviews by SOC practitioners, and codes were finalized through these iterations. We also note that no new code was obtained after the 14th interview, and hence we judged the saturation of themes. Therefore, we did not conduct additional interviews beyond the 15 participants. In the following, we describe the evaluation criteria for assessing report quality that were derived from the interview results.
@@ 3.3.2 Implicit Evaluation Criteria for Security Report Quality
We identify the implicit evaluation criteria that SOC practitioners use in evaluating analysis reports and refine them into explicit evaluation criteria. Although most of the organizations we interviewed did not maintain formal documents of evaluation criteria for analysis reports, we found that several veteran SOC practitioners, such as Tier-2–level analysts and recipient-side managers, have similar criteria when they evaluate analysis reports. Moreover, most comments pointed out the quality of the report content instead of its formal quality. We present the resulting evaluation criteria and their related comments in accordance with the three categories, including the items that should be included in the report.
@@
Criterion (1): Clarity and persuasiveness of the bottom-line conclusion for rapid situation understanding. The participants emphasized that analysis reports should present a clear and persuasive conclusion that enables readers to take and decide next actions. In particular, analysis reports are expected to explicitly document whether security alerts are valid and urgent, the recommended next actions against the alerts, and the reasons that support these judgments. However, the participants noted that many analysis reports merely enumerate factual descriptions without documenting the above information. “High-quality reports clearly state whether an alert should be treated as a true positive or not, how urgent it is, and why that judgment was made. Without that bottom line, it is hard to decide the next step quickly.” –- UserID 1. “For decision-making, the key is whether the report clearly explains what the problem is, why it matters, and what conclusion we should draw from the analysis.” –- UserID 3.
@@
Criterion (2): Quality of concreteness and prioritization in the proposed next actions. The participants noted that analysis reports should not end with analysis results but should clearly outline the next actions with their priorities. Effective analysis reports enable readers to distinguish between actions that must be taken immediately and those that are optional or conditional, thereby allocating resources appropriately. In contrast, analysis reports that simply list insights without actionable guidance are insufficient for operational use. “It is important to clearly state what should be done next and to indicate priorities, for example, whether an action is mandatory or should be handled if resources allow.” –- UserID 9. “Readers need to understand not only what issues exist, but also what actions they themselves are expected to take in response to those issues.” –- UserID 6.
@@
Criterion (3): Adequacy and clarity of the described customer or business impact. The participants regarded it as essential that analysis reports clearly describe how each security alert affects the customer’s business and operations from the customer’s perspective. Analysis reports that focus solely on technical details without considering business impact are often difficult for readers to judge whether the underlying security alerts are acceptable. Quantifying the impact is also valuable for decision-making. “If the impact on business or operations is clearly described or even roughly quantified, it becomes much easier to explain the situation and move forward.” –- UserID 11. “Ultimately, customers want to know what caused the issue and whether the situation is truly acceptable. If that is unclear, the report leaves them uneasy.” –- UserID 14.
@@
Criterion (4): Appropriateness of tailoring the content, tone, and level of detail to the recipient’s role and level of understanding. The participants highlighted that it is important to adjust the content, including its technical details, to readers’ roles and expertise in order to reduce cognitive cost. It also enables the readers to understand whether security alerts are urgent. The participants noted that reports written in a uniform style often fail to bridge the gap between frontline analysts and stakeholders who make decisions, leading to misunderstandings and delayed responses. “Readability is critical. Each recipient has a different level of understanding, so the explanation and terminology need to be adjusted accordingly.” –- UserID 2. “What matters most is narrowing the gap in sense of urgency between the frontline and decision-makers. Reports should reflect the context and discussions that led up to the analysis.” –- UserID 10.
@@
Evaluation criterion (5): Accuracy and explanatory strength of factual descriptions, including explicit assumptions and constraints. The participants emphasized that analysis reports must precisely describe what occurred during the event and provide explanations to convincingly justify the conclusion. The participants highlighted that explicitly stating the assumptions, analytical grounds, and constraints underlying the analysis is important beyond factual descriptions. Analysis reports that lack these contents are often difficult for readers to trust the conclusion. “The most important thing is to accurately describe what actually happened and to explain it in a way that the reader can reasonably accept.” –- UserID 7. “If the assumptions, grounds, or constraints behind the analysis are unclear, it becomes hard to judge whether the situation has really been understood correctly.” –- UserID 4.
@@
Evaluation criterion (6): Appropriateness and depth of on-site contextual information usage, such as asset roles and operational background. The participants noted that proper interpretation of security alerts needs operational context, including asset roles and their position within operational processes. Unless such context is present, analysts are often unable to determine whether each security alert is a genuine issue or a false positive. The participants also noted that insufficient operational context frequently leads to overly ambiguous conclusions. “Operational context is indispensable. Depending on the role of the asset and how it is used, the priority and meaning of the alert can change completely.” –- UserID 13. “If information such as asset roles, whether the traffic is business-related, or results of on-site interviews is missing, we often cannot clearly determine whether an alert is a false positive.” –- UserID 5.
@@
Evaluation criterion (7): Adequacy of technical depth in describing the event beyond simple enumeration of logs. Rather than merely listing communication records and log entries, analysis reports should provide technically meaningful explanations of the event, such as what they are used for. Several participants expressed frustration that interpreting raw technical data is often insufficient. “SOC reports often just list IP addresses that communicated, and we have to re-investigate where those systems are and what they are used for.” –- UserID 6. “If the report also explained the meaning of the communication, it would significantly reduce the investigation workload on our side.” –- UserID 11.
@@
Evaluation criterion (8): Robustness of anomaly identification through multiple analytical approaches, such as trends and comparisons. The participants underscored the importance of identifying anomalies using multiple analytical perspectives, including temporal trends, comparisons with historical data, and correlations across different logs. It is considered that analysis reports that rely solely on isolated log entries are insufficient for understanding whether observed behavior is truly abnormal. The participants noted that temporal and comparative analyses are still underutilized in many current analysis reports. “What we really want to know is what has changed compared to the past, but that perspective is often missing in reports.” –- UserID 9. “Showing insights derived from differences in trends or combinations of logs would make anomalies much clearer.” –- UserID 3.
@@
Evaluation criterion (10): Clarity and verifiability of the analysis process through explicit and coherent links between evidence and conclusions. It was considered critical that reports clearly document the reasoning process by explicitly linking evidence to conclusions, so that a third party could trace and verify how judgments were reached. The participants pointed out that analysis reports often fail to sufficiently document how alternative explanations, such as false positives, were examined, thereby undermining confidence in the conclusions and the overall quality of the analysis. “It is important that there are no logical gaps or contradictions in how the analysis results lead to the conclusions, and that a third party could reproduce the same reasoning.” –- UserID 7. “Especially for determining whether an alert is a false positive, the evidence and reasoning must be clearly connected. If that part is weak, the quality of the entire report appears low.” –- UserID 11.
@@ 3.4 Development of the Analyst-wise Checklist Draft
The Analyst-wise Checklist was developed by systematically integrating two sources, i.e., 1) the structural elements that should be included in analysis reports, which were extracted through a literature review, and 2) the implicit evaluation criteria used in practice by SOC analysts, which were obtained through interviews. The integrated item set was then finalized through expert review and application tests. This approach is also widely known as a general method for constructing evaluation checklists [hales2008development].
@@
When designing checklist items, our primary requirement is that they remain actionable in day-to-day SOC operations. Simply turning participants’ concerns into items risks creating demands that exceed the authority or resources of a typical SOC organization[rose2022co]. We therefore introduced a step that distills practically feasible elements from the interview findings and refines them into checklist items. For example, for evaluation criterion (6): use of on-site contextual information, the participants repeatedly reported that the lack of local context often prevents them from deciding whether an alert is a false positive. However, it is unrealistic for a standard SOC to maintain detailed, up-to-date business context for every customer environment. Accordingly, our checklist does not merely ask whether incidents are analyzed with customer context in mind, but also whether analysts explicitly state which contextual information was available and organize missing yet decision-critical information as concrete confirmation requests. This yields a realistic checklist that assumes a division of roles with on-site staff while still reducing ambiguity caused by insufficient context.
The Analyst-wise Checklist is also applicable to general tasks by human in SOCs because of the following concept: 1) the granularity of checklist items are generalized for the use in multiple domains, 2) for operational difficulty and flexibility in real-world systems, we do not define detailed, item-level scoring rubrics for each checklist item, 3) instead of applying the entire checklist, only relevant items for each analysis report are chosen.
@@ 3.5 Refinement through Expert Review and Evaluation Tests
As a final validation step, one of the three experts and one of the authors with more than seven years of SOC-related experience independently applied the checklist to ten real-world security alert analysis reports. For each checklist item, they rated the extent to which it was satisfied on a five-point scale and recorded the results. For items where their ratings diverged substantially in direction, e.g., one judged the item to be largely satisfied while the other judged it to be not satisfied, they reconciled their interpretations and refined the item descriptions and wording accordingly.
We confirmed that for more than 60% of the checklist items, the two raters agreed on the direction of judgment, i.e., whether the item was satisfied or not, and we then finalized the checklist. Table 4 shows a subset of the checklist with mappings to the evaluation criteria. The complete checklist and checklist item types are provided in Table 5 and Table 6.